OpenStudy (anonymous):

An enclosed gas exerts a pressure P on the walls of a container. This pressure is directly proportional to the temperature T of the gas, measured in Kelvin. If the pressure is 7 lb per square inch when the temperature is 420ºK, find the constant of variation. 49 60 1/49 1/60

OpenStudy (agl202):

P = (k)T => 7 = (k)420 => k = 7/420 = 1/60 = D.

OpenStudy (anonymous):

thank you:)

OpenStudy (agl202):

Anytime!